# KiDS-SBI: Simulation-Based Inference Analysis of KiDS-1000 Cosmic Shear

### v1

@@
(Received xx; accepted xx)
We present a simulation-based inference (SBI) cosmological analysis of cosmic shear two-point statistics from the fourth weak gravitational lensing data release of the ESO Kilo-Degree Survey (KiDS-1000). KiDS-SBI efficiently performs non-Limber projection of the matter power spectrum via Levin’s method, and constructs log-normal random matter fields on the curved sky for arbitrary cosmologies, including effective prescriptions for intrinsic alignments and baryonic feedback. The forward model samples realistic galaxy positions and shapes based on the observational characteristics of KiDS-1000, incorporating shear measurement and redshift calibration uncertainties, as well as angular anisotropies due to variable survey depth and point-spread function variations. To enable direct comparison with standard inference, we limit our analysis to pseudo-angular power spectra as summary statistics. The SBI is based on neural density estimation of the likelihood with active learning to infer the posterior distribution of spatially-flat $\Lambda$CDM cosmological parameters from 18,000 realisations. We infer a mean marginal of the growth of structure parameter $S_{8}\equiv\sigma_{8}(\Omega_{\mathrm{m}}/0.3)^{0.5}=0.731\pm 0.033$ ($68\%$). We present a measure of goodness-of-fit for SBI and determine that the forward model fits the data well with a probability-to-exceed of $0.42$. For fixed cosmology, the learnt likelihood is approximately Gaussian, while constraints widen compared to a Gaussian likelihood analysis due to cosmology dependence in the covariance. Neglecting variable depth and anisotropies in the point spread function in the model can cause $S_{8}$ to be overestimated by ${\sim}5\%$. Our results are in agreement with previous analysis of KiDS-1000 and reinforce a $2.9\sigma$ tension with early-Universe constraints from cosmic microwave background measurements. This work highlights the importance of forward-modelling systematic effects in upcoming galaxy surveys, such as Euclid, Rubin and Roman.
@@ 1 Introduction
Cosmic shear, the weak gravitational lensing effect on distant galaxies due to matter in the foreground, is a powerful tool to study the distribution of matter in the Universe and to probe its large-scale structure. By measuring the distortions in the shapes of galaxy images caused by the gravitational influence of intervening matter, we can infer a combination of the matter density and the amplitude of the matter power spectrum within the framework of a spatially flat cold dark matter cosmology with a cosmological constant, $\Lambda$CDM. Recent cosmic shear analyses of the data taken in Stage-III galaxy surveys, such as the Kilo-Degree Survey [^1] https://kids.strw.leidenuniv.nl/ (KiDS; Kuijken et al. 2019; Asgari et al. 2021; Heymans et al. 2021; Loureiro et al. 2022; van den Busch et al. 2022; Li et al. 2023a), the Hyper Suprime-Cam survey [^2] https://hsc.mtk.nao.ac.jp/ssp/ (HSC; Sugiyama et al. 2022; Aihara et al. 2022; Li et al. 2023b), and the Dark Energy Survey [^3] https://www.darkenergysurvey.org/ (DES; Gatti et al. 2021; Amon et al. 2022; Secco et al. 2022), have constrained these cosmological parameters with unprecedented precision. Upcoming Stage-IV galaxy surveys, such as Euclid [^4] https://www.euclid-ec.org/ (Laureijs et al., 2011), Rubin  [^5] https://www.lsst.org/ (LSST Science Collaboration et al., 2009) or Roman [^6] https://roman.gsfc.nasa.gov/ (Spergel et al., 2015), will further improve upon these constraints, as these surveys will be wider, deeper and more precise. However, many analysis challenges still remain.
@@
We aim to shed some light onto this by conducting a full cosmic shear analysis using neural likelihood estimation simulation-based inference (SBI) with the same complexity as current Stage-III analyses. SBI, also known as Likelihood-Free Inference or Implicit Likelihood Inference, is a Bayesian inference method which does not require an explicit formulation for the likelihood function of the data given the parameters of interest. Instead, the likelihood is implicitly calculated by evaluating the joint probability of data and parameters of interest from forward simulations which map the parameters to the corresponding mock data vectors. This comes with multiple advantages with respect to other standard approaches which require an explicit form for the likelihood. Firstly, the likelihood is allowed to take an arbitrary form, so one can avoid the typical assumption of a Gaussian likelihood or avoid having to define a complex analytical expression for the likelihood. Secondly, for some models and measurements, it may not even be possible to define an analytical likelihood or it would be too resource-intensive, e.g. the covariance of nth-order statistics depends on up to 2nth-order correlation functions which become geometrically more expensive to compute. In such cases, as long as the observables can be simulated, an effective/implicit likelihood may be found using SBI. Lastly, in standard analyses, it often is required to validate the signal and noise modelling with forward simulations, so one can make use of existing tools to perform SBI.
@@
When running a Markov-Chain Monte Carlo (MCMC) chain to sample the posterior distributions from a Gaussian likelihood, one either has to compute a numerical covariance matrix from forward simulations or create an analytical model for the covariance which ought to be validated with a numerical one. In both cases, for a data vector, $\mathbf{d}$, one would require a number of simulations $>|\mathbf{d}|$ to get an accurate sample covariance matrix. This is already the approximate amount of forward simulations needed for an SBI analysis (Alsing et al., 2018; Lin et al., 2023). Therefore, SBI allows one to do full Bayesian uncertainty propagation from data to parameters for any model which can be simulated, without substantial additional computational costs when compared to standard approaches at comparable accuracy.
@@
In cosmic shear, we know that the commonly used two-point statistics to quantify the shear-shear correlations of galaxies on the sky have an approximately Gaussian likelihood, but this assumption has its limitations (Schneider & Hartlap, 2009; Sellentin & Heavens, 2018; Sellentin et al., 2018; Taylor et al., 2019; Upham et al., 2021; Hall & Taylor, 2022). Additionally, when observational and physical systematics are included, the likelihood may become even less Gaussian (Jeffrey et al., 2021). In such cases, like the one presented in this paper, SBI allows one to capture potential non-Gaussanities and propagate them to the posterior distributions.
@@
SBI can also compute effective likelihoods for summary statistics for which analytical methods may be intractable, such as higher-order cosmic shear statistics (Fluri et al., 2022; Euclid Collaboration et al., 2023; Lu et al., 2023). In this work, we limit the scope to only angular power spectra of cosmic shear and leave higher-order statistics or field-level statistics within SBI as an avenue for future work. This allows us to make a direct comparison to previous standard cosmic shear analyses (e.g. Asgari et al. 2021; Loureiro et al. 2022), while also harnessing the efficiency of log-normal forward models of matter density fields which are mostly accurate for two-point statistics (Hall & Taylor, 2022; Piras et al., 2023).
@@
A potential cause for deviations from Gaussianity in the likelihood of two-point statistics/angular power spectra is the anisotropy in the observational depth of a galaxy survey across the sky (Guzik & Bernstein, 2005; Shirasaki et al., 2019). As a survey such as KiDS images the sky, each pointing and each pixel are subject to variable environmental/background conditions which may affect the selection of galaxies (e.g. atmospheric conditions, thermal expansion of the telescope, moonlight, zodiacal light, Galactic extinction, etc.).
@@
Due to the given survey strategy and/or dithering to compensate for gaps in the telescope’s charge-coupled device (CCD), it is possible that there is spatial variation in the overlap between pointings. When the images of each pointing are subsequently combined, some galaxies are observed more often than others which can cause variations in the signal-to-noise ratio which are spatially correlated.
@@
In effect, all these variations lead to fluctuations in the effective galaxy density which are observed across the footprint, while also selecting galaxies according to their shape. This has been found to be a percent-level effect on the cosmic shear signal measured by KiDS (Heydenreich et al., 2020), while also affecting the standard deviation of the KiDS-1000 cosmic shear signal (by $\lesssim 20\%$; Joachimi et al. 2021). Although for KiDS-1000 this is within the statistical uncertainty, it will become important for upcoming galaxy shear surveys such as Euclid, Rubin or Roman. Additionally, the effect of spatial variability on the non-Gaussianity of the likelihood has not been explored in previous work. To address these needs, in this work, our forward model includes the effects of the observational depth variations in KiDS, so that we may propagate the effects of spatial variations in the depth to the uncertainty of inferred cosmological parameters.
@@
To conduct the SBI, we choose to make use of Density Estimation Likelihood-Free Inference (DELFI, Alsing et al. 2019), as it offers good performance and scales better when increasing the dimensionality of the parameter space of interest than other SBI methods (Leclercq, 2018; Alsing et al., 2018), such as Approximate Bayesian Computation (ABC; Rubin 1984; Pritchard et al. 1999; Lin & Kilbinger 2015; Lin et al. 2016; Beaumont 2019). DELFI discards less information from the forward simulations than ABC. ABC only keeps the forward simulations which have a certain level of agreement with the desired data vector. In contrast, DELFI learns a probability density distribution of the data vectors as a function of the model parameters based on all the forward simulations. To achieve this, DELFI employs ensembles of neural density estimators in order to learn the sampling distribution of the data and the parameters from the forward simulations.
@@
Of course, as with any inference method, one needs to be mindful of any intrinsic biases which might come with SBI, how to mitigate them and avoid them. As the likelihood is not known analytically, it is important that the chosen SBI method, sampling scheme and number of simulations are sufficient to ensure that the learnt likelihood has converged to the model’s true likelihood function. To avoid biasing the inference, this requires careful robustness and convergence tests of the implicit likelihood which may not always be necessary for a standard Gaussian likelihood analysis. Hence, we dedicate a considerable effort in this work to ensure the inference pipeline is validated, the neural density estimators of DELFI have converged, the inferred posterior is accurate, and the inferred parameters give a good fit to the data.
@@
SBI is increasing in popularity, among other reasons, because it requires about an order of magnitude fewer evaluations of the likelihood than a standard Gaussian likelihood MCMC analysis to constrain the posterior (Papamakarios & Murray, 2016; Alsing et al., 2018; Gerardi et al., 2021; Jeffrey et al., 2021; Lemos et al., 2021; Legin et al., 2021; Mishra-Sharma & Cranmer, 2022; Spurio Mancini et al., 2023; Lin et al., 2023; Chen et al., 2023; Hu et al., 2022; Lemos et al., 2023b, c; Gatti et al., 2024; Moser et al., 2024). We highlight the methodology of Gatti et al. (2024) which conducted a similar analysis to the one outlined in this paper with the DES Year 3 weak lensing data; see Jeffrey et al. (2024) for results. Whereas the DES analysis focuses on SBI of two-point and higher-order shear statistics, as well as cosmologies beyond $\Lambda$CDM based on a fixed number of N-body simulations while incorporating standard systematics, the analysis in this paper exclusively considers two-point statistics within the context of $\Lambda$CDM. We highlight that in our analysis the data’s likelihood is learned by actively sampling as many efficient forward simulations as needed which capture many physical and observational systematics, including anisotropic systematics which are typically not considered in weak lensing analyses. With this, we aim to give more insight into the robustness against systematics of standard weak lensing analyses.
@@
In a previous paper (Lin et al., 2023, L23 hereafter), we introduced a new inference pipeline for estimating the cosmological parameters from the cosmic shear data of KiDS-1000 (Kuijken et al., 2019) using DELFI. In L23, the SBI pipeline for KiDS is shown to be robust, accurate and efficient, even when constraining a 12-dimensional posterior distribution. However, in that work, the simulated vectors were based on random samples from the covariance matrix used in the fiducial Gaussian KiDS-1000 analysis, so the likelihood was Gaussian by construction to allow for direct comparison with the standard formalism. In this paper, we present a novel suite of physically motivated and realistic forward simulations of the cosmic shear observables as seen by KiDS, so the neural density estimators learn any non-Gaussianities in the likelihood induced by our model, plus any complex relations between the data and the model parameters. We then use these simulations to create two inference pipelines: a fiducial one which makes the same modelling choices as the KiDS-1000 analysis (Joachimi et al., 2021) and a more realistic one which explicitly models systematic effects that would be prohibitively difficult to incorporate into a standard likelihood analysis. From each pipeline, we obtain a separate posterior estimate for $S_{8}$ which we then scrutinise with a novel measure of goodness-of-fit for SBI posteriors.
@@
This paper has the following structure: Sect. 2 presents the KiDS data used for this analysis. Section 3 gives a detailed description of the forward simulations developed for the SBI analysis. Section 4 details how the galaxy shapes and summary statistics are measured in both the simulations and the data. Section 5 describes the setup of the simulation-based inference pipeline using DELFI based on L23. Section 6 presents the novel measure of goodness-of-fit for SBI that we use in this paper. In Sect. 7, we conduct consistency tests to ensure the SBI pipeline is unbiased and robust, while Sect. 8 discusses the validity of a Gaussian likelihood assumption. Section 9 presents the cosmological inference results from the KiDS-1000 weak lensing data. Lastly, we give concluding remarks in Sect. 10, and provide additional details about the analysis in the Appendix.
Refer to captionRefer to caption
@@
Figure 2: The redshift distributions of the five KiDS-1000 tomographic bins. The shaded areas show to limits of each tomographic bin, while the solid lines show the $n(z)$ of the source galaxies in each tomographic bin as a function of both redshift, $z$, and comoving distance, $\chi$ (the latter is derived assuming a Planck 2018 cosmology; Planck Collaboration et al. 2020). The black dashed lines show the limits of the spherical matter shells in our forward simulations.

### v2

@@
(Received 23 April 2024 / Accepted 10 December 2024)
We present a simulation-based inference (SBI) cosmological analysis of cosmic shear two-point statistics from the fourth weak gravitational lensing data release of the ESO Kilo-Degree Survey (KiDS-1000). KiDS-SBI efficiently performs non-Limber projection of the matter power spectrum via Levin’s method and constructs log-normal random matter fields on the curved sky for arbitrary cosmologies, including effective prescriptions for intrinsic alignments and baryonic feedback. The forward model samples realistic galaxy positions and shapes, based on the observational characteristics of KiDS-1000. It incorporates shear measurement and redshift calibration uncertainties, as well as angular anisotropies due to variable survey depth and point spread function (PSF) variations. To enable direct comparisons with standard inference, we limited our analysis to pseudo-angular power spectra as summary statistics. Here, the SBI is based on neural density estimation of the likelihood with active learning to infer the posterior distribution of spatially flat $\Lambda$CDM cosmological parameters from 18,000 realisations. We inferred a mean marginal for the growth of the structure parameter of $S_{8}\equiv\sigma_{8}(\Omega_{\mathrm{m}}/0.3)^{0.5}=0.731\pm 0.033$ ($68\%$). We present a measurement of the goodness-of-fit for SBI, determining that the forward model fits the data well, with a probability-to-exceed of $0.42$. For a fixed cosmology, the learnt likelihood is approximately Gaussian, while its constraints are wider, compared to a Gaussian likelihood analysis due to the cosmology dependence in the covariance. Neglecting variable depth and anisotropies in the point spread function in the model can cause $S_{8}$ to be overestimated by ${\sim}5\%$. Our results are in agreement with previous analyses of KiDS-1000 and reinforce a $2.9\sigma$ tension with early Universe constraints from cosmic microwave background measurements. This work highlights the importance of forward-modelling systematic effects in upcoming galaxy surveys, such as Euclid, Rubin, and Roman.
@@ 1 Introduction
The weak gravitational lensing effect on distant galaxies due to matter in the foreground, known as cosmic shear, is a powerful tool for studying the distribution of matter in the Universe and to probe its large-scale structure. By measuring the distortions in the shapes of galaxy images caused by the gravitational influence of intervening matter, we can infer a combination of the matter density and the amplitude of the matter power spectrum within the framework of a spatially flat cold dark matter cosmology with a cosmological constant, $\Lambda$CDM. Recent cosmic shear analyses of the data taken in stage-III galaxy surveys, such as the Kilo-Degree Survey [^1] https://kids.strw.leidenuniv.nl/ (KiDS; Kuijken et al. 2019; Asgari et al. 2021; Heymans et al. 2021; Loureiro et al. 2022; van den Busch et al. 2022; Li et al. 2023a), the Hyper Suprime-Cam survey [^2] https://hsc.mtk.nao.ac.jp/ssp/ (HSC; Sugiyama et al. 2022; Aihara et al. 2022; Li et al. 2023b), and the Dark Energy Survey [^3] https://www.darkenergysurvey.org/ (DES; Gatti et al. 2021; Amon et al. 2022; Secco et al. 2022), have constrained these cosmological parameters with unprecedented precision. Upcoming stage-IV galaxy surveys, such as Euclid [^4] https://www.euclid-ec.org/ (Laureijs et al., 2011), Rubin [^5] https://www.lsst.org/ (LSST Science Collaboration et al., 2009), and Roman [^6] https://roman.gsfc.nasa.gov/ (Spergel et al., 2015), will further improve upon these constraints, as these surveys will be wider, deeper, and more precise. However, many analytical challenges still remain.
@@
We aim to shed some light on this issue by conducting a full cosmic shear analysis using neural likelihood estimation simulation-based inference (SBI) characterised by the same complexity as current stage-III analyses. SBI, also known as ’likelihood-free inference’ or implicit likelihood inference’, is a Bayesian inference method that does not require an explicit formulation for the likelihood function of the data given the parameters of interest. Instead, the likelihood is implicitly calculated by evaluating the joint probability of data and parameters of interest from forward simulations, which map the parameters to the corresponding mock data vectors. This comes with multiple advantages with respect to other standard approaches that require an explicit form for the likelihood. Firstly, the likelihood is allowed to take an arbitrary form, so we can avoid the typical assumption of a Gaussian likelihood or avoid having to define a complex analytical expression for the likelihood. Secondly, for some models and measurements, it may not even be possible to define an analytical likelihood or it would be too resource-intensive; for instance, the case of covariance among nth-order statistics depending on up to 2nth-order correlation functions, which become geometrically more expensive to compute. In such cases, as long as the observables can be simulated, an effective or implicit likelihood may be found using SBI. Lastly, in standard analyses, it often is required to validate the signal and noise modelling with forward simulations, so that we can make use of existing tools to perform SBI.
@@
When running a Markov chain Monte Carlo (MCMC) to sample the posterior distributions from a Gaussian likelihood, we either have to compute a numerical covariance matrix from forward simulations or create an analytical model for the covariance, which ought to be validated with a numerical one. In both cases, for a data vector, $\mathbf{d}$, we would require a number of simulations $>|\mathbf{d}|$ to get an accurate sample covariance matrix. This is already the approximate amount of forward simulations needed for an SBI analysis (Alsing et al., 2018; Lin et al., 2023). Therefore, SBI allows us to perform a full Bayesian uncertainty propagation from the data to obtain parameters for any model that can be simulated, without substantial additional computational costs when compared to standard approaches at a comparable accuracy.
@@
With respect to cosmic shear, we know that the commonly used two-point statistics to quantify the shear-shear correlations of galaxies on the sky have an approximately Gaussian likelihood, but this assumption has its limitations (Schneider & Hartlap, 2009; Sellentin & Heavens, 2018; Sellentin et al., 2018; Taylor et al., 2019; Upham et al., 2021; Hall & Taylor, 2022). Additionally, when observational and physical systematics are included, the likelihood may become even less Gaussian (Jeffrey et al., 2021). In such cases, like the one presented in this paper, SBI allows us to capture potential non-Gaussanities and propagate them to the posterior distributions.
@@
The SBI approach can also enable us to compute effective likelihoods for summary statistics for which analytical methods may be intractable, such as higher order cosmic shear statistics (Fluri et al., 2022; Euclid Collaboration et al., 2023; Lu et al., 2023). In this work, we limit the scope to only angular power spectra of cosmic shear and leave higher order statistics or field-level statistics within SBI as an avenue for future work. This allows us to make a direct comparison to previous standard cosmic shear analyses (such as Asgari et al. 2021; Loureiro et al. 2022), while also harnessing the efficiency of log-normal forward models of matter density fields which are mostly accurate for two-point statistics (Hall & Taylor, 2022; Piras et al., 2023).
@@
A potential cause for deviations from Gaussianity in the likelihood of two-point statistics/angular power spectra is the anisotropy in the observational depth of a galaxy survey across the sky (Guzik & Bernstein, 2005; Shirasaki et al., 2019). As a survey such as KiDS images the sky, each pointing and each pixel are subject to variable environmental and background conditions that may affect the selection of galaxies (for example atmospheric conditions, thermal expansion of the telescope, moonlight, zodiacal light, Galactic extinction, etc.).
@@
Due to the given survey strategy and/or dithering to compensate for gaps in the telescope’s charge-coupled device (CCD), it is possible that there is spatial variation in the overlap between pointings. When the images of each pointing are subsequently combined, some galaxies are observed more often than others that can cause variations in the signal-to-noise ratio (S/N), which are spatially correlated.
@@
In effect, all these variations lead to fluctuations in the effective galaxy density, which are observed across the footprint, while also selecting galaxies according to their shape. This has been found to be a percent-level effect on the cosmic shear signal measured by KiDS (Heydenreich et al., 2020), while also affecting the standard deviation of the KiDS-1000 cosmic shear signal (by $\lesssim 20\%$; Joachimi et al. 2021). For KiDS-1000, this is within the statistical uncertainty, however, it will become important for upcoming galaxy shear surveys such as Euclid, Rubin, and Roman. Additionally, the effect of spatial variability on the non-Gaussianity of the likelihood has not been explored in previous works. To address these needs, in this work, our forward model includes the effects of the observational depth variations in KiDS, so that we may propagate the effects of spatial variations in the depth to the uncertainty of inferred cosmological parameters.
@@
To conduct the SBI, we chose to make use of the density estimation likelihood-free inference (DELFI, Alsing et al. 2019) method, as it offers good performance and scales better when increasing the dimensionality of the parameter space of interest than other SBI methods (Leclercq, 2018; Alsing et al., 2018), such as approximate Bayesian computation (ABC; Rubin 1984; Pritchard et al. 1999; Lin & Kilbinger 2015; Lin et al. 2016; Beaumont 2019). DELFI discards less information from the forward simulations than ABC, whereas the latter only keeps the forward simulations which have a certain level of agreement with the desired data vector. In contrast, DELFI learns a probability density distribution of the data vectors as a function of the model parameters based on all the forward simulations. To achieve this, DELFI employs ensembles of neural density estimators to learn the sampling distribution of the data and the parameters from the forward simulations.
@@
Of course, as with any inference method, we need to be mindful of any intrinsic biases that might come with SBI, as well as how to mitigate and avoid them. As the likelihood is not known analytically, it is important that the chosen SBI method, sampling scheme, and number of simulations are sufficient to ensure that the learnt likelihood has converged to the model’s true likelihood function. To avoid biasing the inference, this requires careful robustness and convergence tests of the implicit likelihood, which may not always be necessary for a standard Gaussian likelihood analysis. Hence, we dedicate a considerable effort in this work to ensure the inference pipeline is validated, the neural density estimators of DELFI have converged, the inferred posterior is accurate, and the inferred parameters give a good fit to the data.
@@
Applications of SBI are increasing in popularity, in major part because it requires about an order of magnitude fewer evaluations of the likelihood than a standard Gaussian likelihood MCMC analysis to constrain the posterior (Papamakarios & Murray, 2016; Alsing et al., 2018; Gerardi et al., 2021; Jeffrey et al., 2021; Lemos et al., 2021; Legin et al., 2021; Mishra-Sharma & Cranmer, 2022; Spurio Mancini et al., 2023; Lin et al., 2023; Chen et al., 2023; Hu et al., 2022; Lemos et al., 2023b, c; Gatti et al., 2024; Moser et al., 2024; Franco-Abellán et al., 2024). We highlight the methodology of Gatti et al. (2024), which used a similar analysis to the one outlined in this paper, with the DES Year 3 weak lensing data. We refer to Jeffrey et al. (2024) for a full description of the results. While the DES analysis focuses on SBI of two-point and higher-order shear statistics, along with cosmologies beyond $\Lambda$CDM based on a fixed number of N-body simulations, while incorporating standard systematics, the analysis in this paper exclusively considers two-point statistics within the context of $\Lambda$CDM. We highlight that in our analysis, the data likelihood is learned by actively sampling as many efficient forward simulations as needed. This allows us to capture many physical and observational systematics, including anisotropic systematics that are typically not considered in weak lensing analyses. With this, we aim to give more insight into the robustness against systematics of standard weak lensing analyses.
@@
In a previous paper (Lin et al., 2023, L23 hereafter), we introduced a new inference pipeline for estimating the cosmological parameters from the cosmic shear data of KiDS-1000 (Kuijken et al., 2019) using DELFI. In L23, the SBI pipeline for KiDS is shown to be robust, accurate and efficient, even when constraining a 12-dimensional (12D) posterior distribution. However, in that work, the simulated vectors were based on random samples from the covariance matrix used in the fiducial Gaussian KiDS-1000 analysis, so the likelihood was Gaussian by construction to allow for direct comparison with the standard formalism. In this paper, we present a novel suite of physically motivated and realistic forward simulations of the cosmic shear observables, as seen by KiDS. Thus, the neural density estimators can learn any non-Gaussianities in the likelihood induced by our model, plus any complex relations between the data and the model parameters. We then used these simulations to create two inference pipelines: a fiducial one which makes the same modelling choices as the KiDS-1000 analysis (Joachimi et al., 2021) and a more realistic one, which explicitly models systematic effects that would be prohibitively difficult to incorporate into a standard likelihood analysis. From each pipeline, we obtained a separate posterior estimate for $S_{8}$, which we then scrutinised with a novel measure of the goodness-of-fit for SBI posteriors.
@@
This paper has the following structure. Section 2 presents the KiDS data used for this analysis. Section 3 gives a detailed description of the forward simulations developed for the SBI analysis. Section 4 details how the galaxy shapes and summary statistics are measured in both the simulations and the data. Section 5 describes the setup of the simulation-based inference pipeline using DELFI based on L23. Section 6 presents the novel measure of goodness-of-fit for SBI we use in this paper. In Sect. 7, we describe our consistency tests to ensure the SBI pipeline is unbiased and robust, while in Sect. 8, we discusse the validity of a Gaussian likelihood assumption. Section 9 presents the cosmological inference results from the KiDS-1000 weak lensing data. Lastly, we give our concluding remarks in Sect. 10. We provide additional details about the analysis in the appendix.
@@
Figure 2: Redshift distributions of the five KiDS-1000 tomographic bins. The shaded areas show to limits of each tomographic bin, while the solid lines show the $n(z)$ of the source galaxies in each tomographic bin as a function of both redshift, $z$, and comoving distance, $\chi$ (the latter is derived assuming a Planck 2018 cosmology; Planck Collaboration et al. 2020). The black dashed lines show the limits of the spherical matter shells in our forward simulations.